OpenStudy (anonymous):

the answer to this is supposed to be 2 but I cant get there -2(x-4)-(8x-3)=2-(6x-1)can anyone help me with this

OpenStudy (tkhunny):

-2(x-4)-(8x-3)=2-(6x-1) What did you do first?

OpenStudy (anonymous):

distribute the -2(x-4)

HanAkoSolo (jamierox4ev3r):

I will do my best -2(x-4)-(8x-3)=2-(-6x-1). I think this is the problem if i understood correctly right? @michaelhale

OpenStudy (anonymous):

yes that is the problem

OpenStudy (tkhunny):

Don't just tell me - SHOW me!

OpenStudy (anonymous):

-2(x-4),-2x+8

HanAkoSolo (jamierox4ev3r):

-2x+8-8x+3=2+6x+1

HanAkoSolo (jamierox4ev3r):

combine the like terms then solve like a regular equation

OpenStudy (anonymous):

how did you get +3

HanAkoSolo (jamierox4ev3r):

a negative (-) times -3 = a positive number, hence +3

HanAkoSolo (jamierox4ev3r):

that is how i distributed it

OpenStudy (tkhunny):

Don't be afraid to rewrite things a bit. It will help you win the battle. -2(x-4)-(8x-3)=2-(6x-1) First parentheses -2x + 8 - (8x-3) = 2 - (6x-1) Next parentheses -2x + 8 - 8x + 3 = 2 - (6x-1) Last parentheses -2x + 8 - 8x + 3 = 2 - 6x + 1 See how easy that is to follow, now that we have written the entire expression every step of the way? Okay, now what?

OpenStudy (anonymous):

so when I see -(8x_3) I distribute the - sign as a coefficient of -1

OpenStudy (tkhunny):

That is a reasonable understanding. It is the same on the last parentheses.

OpenStudy (anonymous):

thank you both I have been beating myself up over this one for a while I got the answer I needed

OpenStudy (tkhunny):

Woo-hoo! Good work!

HanAkoSolo (jamierox4ev3r):

nice process @tkhunny awesome job! ^.^

OpenStudy (anonymous):

i cant give a best response because you both did awsome

OpenStudy (tkhunny):

No worries. You learning is what is important!

OpenStudy (anonymous):

-2x+4-8x+3=2-6x+1 =>-10x+5=-6x+3 =>-4x=-2 =>x=2